# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v1

@@ 2.3.4 Training Procedure
Next, following the assessment of our segmentation results, we introduce example applications to study the utility of the sea, land, and cloud segments in increasing the satellite’s autonomy by decreasing postprocessing demands and reducing decision-making latency. In Fig. 3, we propose a novel sea-land-cloud ranking system for satellite imagery. This system prioritizes HS images for downlink based on sea, land, and cloud coverage levels, ensuring high-priority transmission. The images are arranged in ascending order of cloud coverage levels and descending order of sea and land coverage levels, with the first positions reserved for images of higher priority. Unlike existing ranking systems that target only clouds [12], our approach is the first of its kind in the field, simultaneously identifying and ranking water, land, and clouds within priority queues for HSI-equipped satellites. In Fig. 3, we also suggest other use cases that further highlight the usefulness of the sea-land-cloud segments. These use cases include alternatives such as discarding or applying lossy compression to non-essential pixels at the satellite’s edge to enhance the efficient utilization of the Shannon channel capacity, particularly vital for satellites like HYPSO-1 that may be operating with limited line-of-sight (LOS) towards ground stations [26].
@@
We introduce additional metrics for the ranking system to assist the evaluation of the ML and DL models. To assess the predicted rankings for each coverage level, we use the Spearman’s rank correlation coefficient. This measures the relationship between the predicted and ground-truth rankings, with values ranging from 1-1 to 1, representing a perfect match at the upper end. Moreover, we compute the Mean Absolute Error (MAE) for each coverage level, treating each level as a separate regression task. To maintain simplicity, we include the details about MAE in our supplementary materials.
@@
In the context of on-board data management (also in Fig. 3), the false alarm ratios due to misclassifications indicate either the potential loss of data or the transmission of irrelevant data during downlink, thereby diminishing downlink efficiency. In cloud segmentation, the True Positive Ratio (TPR) involves correctly handling cloudy pixels, while the True Negative Ratio (TNR) relates to recognizing non-cloudy data with potential value. The False Positive Ratio (FPR) is crucial because it indicates misclassified pixels, wrongly identified as clouds but that actually correspond to sea or land. This can result in the elimination or lossy compression of valuable data. On the other hand, the False Negative Ratio (FNR) is less critical than FPR, as it does not involve data loss, but only an inefficient usage of the downlink transmitting undetected pixels actually corresponding to clouds. In the context of sea segmentation, true positives signify correct identification of valuable water data, while true negatives correctly recognize non-relevant data for potential deletion or compression. In sea-focused campaigns, FNR is however critical, as it deals with misclassified sea-related pixels, potentially leading to loss of valuable data. In contrast, FPR is of less significance since its impact lies only in downlink efficiency through the transmission of irrelevant pixels (clouds or land), rather than causing actual data loss. The same analysis and conclusion apply when examining the context of land segmentation. In summary, without taking into account the earlier suggested compromise between false positives and false negatives, for both sea and land categories, FNR has priority over FPR due to a likelihood for valuable data loss in the case of high FNR. However, for cloud data, FPR is critical as it leads to irreversible data loss, with FNR affecting merely downlink efficiency.
@@ 3 Results
We conduct model training using 112 channels of both raw data and L1b radiance in the test set (further discussed in Section 4). Table II presents the outcomes of the segmentation models trained on 112 channels in L1b radiance. In the table, we utilize bold font to accentuate the novel models. The table displays the models’ performance for segmentation, with the primary metric being the average accuracy across the sea, land, and cloud classes. Additionally, to assess the sea-land-cloud segmentation for the context of the on-board ranking system, we also include the average Spearman’s rank coefficient as performance metric. In particular, the mean coefficient across the respective rankings based on sea, land, and cloud coverage levels. Moreover, considering the extensive number of parameters in DL models, which in general can be a challenge for on-board implementations, we also include information about the network’s parameter count to offer insights of the potential memory demands for in-orbit deployment. In addition, we provide the average inference time per HS image measured on ground, on the NVIDIA RTX A4000 GPU (for DL), and on the 13th Gen Intel(R) Core (TM) i9-13900K CPU (for ML). From the numbers presented in Table II, we sort the models in Table III based on their respective metrics. In Table III, we first sort the models solely based on their average accuracy, from highest to lowest, and subsequently, we sort them in the following columns, first by the Spearman’s coefficient (highest to lowest), then by the number of parameters (lowest to highest), and finally by inference time (shorter to longer). Therefore, the first rows in the table represent the top models for each respective metric. However, as the average accuracy only provides a broad overview of the mean segmentation performance across the sea, land, and cloud classes, we additionally provide fine-grained insights into the models’ performance for each separate class by presenting their false detection ratios.
In Figs. 4, 5, and 6, we provide a comparative visual representation of false alarm ratios for the segmentation of each distinct class. We note that the top-performing models get positioned in the lower-left corner, indicating the optimal trade-off balance between false positives and false negatives. Within the context of the example case (Section 2.4), the graphs illustrate the false alarm ratio on the Y-axis corresponding to the loss of valuable data, while the X-axis pertains to a false detection ratio that reflects the inefficient utilization of the downlink. In relation to this trade-off, we stress that while minimizing the loss of valuable data remains crucial, we also highlight the importance of maintaining an acceptable downlink efficiency. Finally, our experiments conclude by retraining the models using the 3-channel dataset. We employ this dataset to assess the impact of the channel reduction (Section 2.2) on the models’ performance, parameter count, and inference times, as shown in Table IV. We note that the table omits the 1D models due to the 3-channel dataset’s restricted spectral information, as mentioned earlier (Section 2.3.4). Notably, the empirical results are further confirmed by retraining once again the models with a different set of three spectral channels. These new results align with the previous findings. For simplicity, we include these new figures in the supplementary material. As previously noted, the dataset predominantly contains instances of sea and land, with clouds representing the minority class. Consequently, the test set also exhibits a lower proportion of cloud data (14.13%), in contrast to sea (39.01%) and land (46.85%). We recognize this as a limitation that demands future refinement. To mitigate the impact of the class imbalance during the testing phase, we conduct additional experiments. Namely, we perform inference on an additional deployment set specifically designed to test the model’s generalization on e.g. cloudy images to further support the findings of our work, and we provide the results in the supplementary material.
@@ 4 Discussion
4.1 Analysis of Results for Models Trained on 112 Spectral Channels
We first analyze the results for the models trained on 112 channels (Tables II and LABEL:Table:models_sorted_according_to_performance). Our objective is to investigate how the utilization of spatial context by 2D models affects the sea-land-cloud segmentation across test images. We assess the models’ strengths and limitations in terms of performance, spanning from the average accuracy and the false alarm ratios to the Spearman’s correlation coefficient for the ranking system. We finalise this analysis by discussing the model’s parameter count and inference time, which are relevant metrics for on-board deployment. Prior to the analysis of all the metrics, we intuitively formulate a hypothesis, independent of the empirical findings. The hypothesis asserts that it is reasonable to anticipate better performance for 2D models compared to 1D models. This expectation arises because 2D models, unlike 1D models relying solely on spectral context, analyse a richer context comprising both spectral and spatial information. We proceed next to elaborate whether the empirical results substantiate such hypothesis.
@@ 4.1.1 Segmentation Performance (I): Average Accuracy
When examining the segmentation performance solely based on the average accuracy across sea, land, and cloud classes, we observe that our proposed model, 1D-Justo-LiuNet, achieves the highest accuracy of 0.93 (Table II). The result of a 1D model outperforming 2D models is not an exception, as four out of five models ranking among the top performers in accuracy in Table LABEL:Table:models_sorted_according_to_performance are consistently 1D models. The only exception in this top five ranking is our 2D model, 2D-Justo-UNet-Simple, which achieves an accuracy of 0.92, closely approaching the performance of the 1D-Justo-LiuNet. These findings lead to the conclusion that the initially expected hypothesis suggesting that 2D models would outperform 1D models is not valid, as the 2D models actually perform worse. This phenomenon can be attributed to the fact that 1D models, unlike 2D using the spatial context, are not affected by the modest spatial resolution of HYPSO-1’s images while not suffering from the spatial variations that may be practically irrelevant for pixel-level segmentation. Instead, 1D models dedicate their efforts to analyzing exclusively the spectral context, which is a valuable domain in HS imagery. Additionally, Table LABEL:Table:models_sorted_according_to_performance demonstrates that our proposed models (1D-Justo-LiuNet, 2D-Justo-UNet-Simple, 1D-Justo-HuNet, and 1D-Justo-LucasCNN) outperform, in average accuracy, the state-of-the-art models under analysis, when trained on the HYPSO-1 dataset for sea-land-cloud segmentation. Indeed, we surpass models such as the 2D-NU-Net-mod, 2D-NU-Net, 2D-UNet FAUBAI, and 2D-UNet FAUBAI Reduced previously employed in the field of cloud segmentation for remote sensing imagery [12, 39]. Finally, we highlight that the average accuracy of 1D-Justo-LiuNet, 1D-Justo-HuNet, and 1D-Justo-LucasCNN (0.93, 0.91, and 0.91, respectively) also surpass the literature’s average accuracy for soil texture. As a matter of fact, according to [42], the state of the art achieves significantly lower average accuracy of 0.59, 0.61, and 0.56 for Liuetal, Huetal, and LucasCNN, respectively.
@@ 4.1.2 Segmentation Performance (II): The Effect of L1b Radiance Calibration
The models in Table II are trained on L1b calibrated radiance data. L1b radiance is one of the standard data formats in HS imagery, which not only allows satellite comparisons but also removes sensor-specific characteristics and biases. Next, we empirically examine how L1b radiance calibration impacts segmentation performance. To be brief while supporting the next discussion, we include the average accuracy for the models trained on raw data in the supplementary materials of this article to allow the comparison with the models trained on L1b radiance in Table II. For the highest-performing DL models in terms of average accuracy shown in Table III (1D-CNNs), they exhibit slightly superior performance with L1b radiance calibration compared to training them with raw data. A similar trend occurs for the 2D models ranked just below the 1D models, which also demonstrate enhanced performance with radiance data versus raw data. Conversely, the models at the lower end of the accuracy ranking tend to perform better with raw data than radiance calibration. Concentrating only on the superior models in Table III, which hold more significance as they are the top performers, we conclude that L1b radiance calibration tends to positively impact segmentation performance. We attribute the improvement to the correction of errors and biases during L1b radiance calibration, allowing top-performing models to focus on inherent data patterns rather than overfitting on sensor-induced characteristics, leading to improved generalization performance. Nonetheless, we strongly recommend conducting comparative testing on both calibrated and raw data when training the models on datasets distinct from this paper to verify whether calibrated data, such as radiance or reflectance, continues to enhance model performance.
@@
4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Efficiency
@@
For a more in-depth assessment of the segmentation performance for each separate class, we scrutinize which models are positioned in the lower-left corner in the scatter plots in Figs. 4, 5, and 6. The bottom-left region ensures the most optimal false detection ratios, and for the example system in Fig. 3, the region achieves minimal loss of valuable data, while also minimizing downlink inefficiency. Next, we comparatively analyze the top-performing models located in the lower-left corner of Figs. 4-6.
As previously explained, lower data loss (in the Y-axis) takes precedence, while the downlink inefficiency (in the X-axis) is of lower priority since it does not result in any loss of valuable data. However, as discussed in Section 2.4, we also aim for a reasonably balanced relationship between false positives and false negatives to prevent scenarios where a model achieves a convenient minimal data loss, at the expense of extremely degrading the downlink efficiency.
@@
First, we assess the models in terms of data loss. In Fig. 4 for cloud segmentation, models such as the 2D-Justo-UNet-Simple excels in minimizing the loss of data, approaching nearly no data loss (ideal). This signifies that the 2D-Justo-UNet-Simple outperforms by not misclassifying valuable non-cloud data (sea or land) as clouds, even if it means some cloudy data goes undetected, reducing the downlink efficiency. With respect to sea segmentation (Fig. 5), the 2D-UNet-FAUBAI-reduced stands out for its minimal loss of sea pixels (FNR=0.01). This indicates that there is only a 1% proportion of sea pixels that remain undetected, potentially resulting in their loss or degradation. This outperforms other models such as the 1D-Justo-LiuNet with 5% loss instead. Finally, in Fig. 6 for land segmentation, we observe a remarkable similarity in false alarm ratios in the lower-left optimal region, resulting in a slight overlap. Therefore, we next clarify that the optimal models, sorted in ascending order of FNR (data loss) values, are as follows: 0.02 (2D-Justo-UNet-Simple), 0.03 (1D-Justo-LiuNet), 0.04 (1D-Justo-HuNet), and 0.06 (2D-CUNet++ Reduced). Consequently, the 2D-Justo-UNet-Simple is the most optimal at minimizing the loss of land data, achieving only a 2% ratio of undetected land pixels.
@@
Second, to jointly maintain a reasonably acceptable downlink efficiency and low data loss, we study the models considering the trade-off between both loss of data and downlink efficiency objectives. One could consider the use of Pareto fronts, which are typically applied in multi-objective optimization to address trade-offs between inversely proportional variables. Nevertheless, for simplicity, we measure the trade-off using the Euclidean distance metric that gives equal weight to the two objectives. Specifically, we calculate the distance of each model from the origin point, representing the ideal scenario with zero false alarms. First, in cloud segmentation (Fig. 4), the models closest to the ideal performance have (FPR, FNR) values of (0.02, 0.15) for 1D-Justo-LiuNet and (0.00, 0.17) for 2D-Justo-UNet-Simple, resulting in approximate distances of 0.151 and 0.170, respectively. In brief, 1D-Justo-LiuNet exhibits the shortest distance to the ideal model. Second, the top models for sea segmentation (Fig. 5), with (FNR, FPR) of (0.08, 0.01) for 1D-Justo-LucasCNN, (0.05, 0.02) for 1D-Justo-LiuNet, and (0.04, 0.03) for 2D-UNet-FAUBAI, have approximate distances of 0.081, 0.054, and 0.050, respectively. This suggests a marginally better performance of 2D-UNet-FAUBAI, with a difference of 0.004 in distance compared to the 1D-Justo-LiuNet. However, Table II notes that the 2D-UNet-FAUBAI has over 26 million parameters, whereas our network, 1D-Justo-LiuNet, has fewer than 5 thousand parameters, which is more suitable for on-board deployment. Considering the nearly identical performance and the large parameter difference against FAUBAI’s network, 1D-Justo-LiuNet remains the top model for sea segmentation. Finally, for land (Fig. 6), the optimal segmentation models are of (FNR, FPR) = (0.02, 0.06) for 2D-Justo-UNet-Simple, (0.03, 0.04) for 1D-Justo-LiuNet, (0.04, 0.04) for 1D-Justo-HuNet, and (0.06, 0.04) for 2D-CUNet++ Reduced, with approximate distances of 0.063, 0.050, 0.057, and 0.072, respectively. In summary, 1D-Justo-LiuNet outperforms the other models for land segmentation.
@@
Considering the initial analysis focused solely on data loss and the subsequent analysis regarding the trade-off between data loss and downlink efficiency, we next provide our conclusion. The 2D-Justo-UNet-Simple, with approximately 7 thousand parameters, minimizes the loss of valuable data for most classes. The only exception is the sea class, where it incurs a 5% data loss, in contrast to the 1% loss by 2D-UNet-FAUBAI-reduced which uses around 2 million parameters and is less suitable for on-board inference. However, when we equally consider both data loss and efficient downlink usage, our model, 1D-Justo-LiuNet, excels in minimizing data loss, achieving rates of 2%, 5%, and 3%, while demonstrating relatively minimal inefficiency rates of 15%, 2%, and 3% for cloud, sea, and land classes, respectively. These findings suggest that when employing 1D-Justo-LiuNet, efficiency slightly deteriorates due to a less precise detection of clouds, leading to an overabundance of cloud data that would be transmitted from the satellite (15%). We attribute this issue to the dataset’s class imbalance, where the cloud class constitutes the minority. Nevertheless, the over-detection of sea (2%) and land (3%) pixels is smaller, thereby maximizing the efficiency of the downlink. Additionally, in our supplementary material, we note that model misclassifications resulting from false alarms predominantly manifest spatially at class edges, where pixels are predicted with considerably lower confidence. This phenomenon is primarily attributed to the inaccuracy of the semi-automatic pixel-level annotation process at class boundaries outlined in our previous publication [31].

### v2

@@ 2.4.1 Proposed On-Board Sea-Land-Cloud Ranking System
Before introducing the metrics to evaluate the segmentation outcomes, we next introduce example applications to study the utility of the sea, land, and cloud segments in increasing the satellite’s autonomy, while reducing data postprocessing and decision-making latency. Several performance metrics discussed in the following section are indeed based on the proposed example applications. Namely, in Fig. 2 we suggest a novel sea-land-cloud ranking system intended for in-orbit deployment. The system prioritizes HS images for downlink based on sea, land, and cloud coverage levels, ensuring high-priority transmission. The images are arranged in ascending order of cloud coverage levels, descending order of sea coverage levels, and also decreasing order of land coverage levels, with the first positions reserved for images of higher downlink priority. Unlike existing ranking systems that target only clouds [13], our proposed approach is the first of its kind for HSI-equipped satellites, simultaneously locating and ranking sea, land, and clouds within priority queues. In Fig. 2, we also suggest other on-board use cases that further highlight the usefulness of the sea-land-cloud segments. The use cases include alternatives such as discarding or applying selective lossy compression to non-essential pixels to optimize the utilization of the Shannon channel capacity, particularly vital for satellites like HYPSO-1 that may be operating with limited line-of-sight (LOS) towards ground stations [29].
Figure 2: An intelligent on-board processing pipeline integrating a ranking system based on sea-land-cloud segmentation. $T\!H_{cl}$, $T\!H_{sea}$, and $T\!H_{land}$ represent thresholds configured at the ground center to regulate data management during in-flight operations.
2.4.2 Performance Metrics
Regarding performance metrics, we utilize a range of metrics to assess the quality of the segmented images on the test set (prior to generalization tests in the deployment set) in order to evaluate the models’ performance after training. First, our primary focus is on average accuracy, a widely used metric for evaluating segmentation models. Unlike overall accuracy, this average metric is particularly convenient for imbalanced test sets, providing a more realistic assessment. Furthermore, adhering to the literature of cloud detection in satellite imagery [12, 14], we also calculate the false and true alarm ratios, for each class in this work. Finally, in our supplementary material supporting further the conclusions of our work, we present additional state-of-the-art metrics that we consider secondary in importance for our purpose compared to the previous metrics. Namely, the F-1 score in addition to the overall accuracy.
@@ 2.4.2 Performance Metrics
Furthermore, we introduce additional performance metrics for the ranking system to assist the evaluation of the ML and DL models. To assess the predicted rankings for each coverage level, we use the Spearman’s rank correlation coefficient to measure the relationship between the predicted and ground-truth rankings, with values ranging from 1-1 to 1, representing a perfect match at the upper end. Moreover, we compute the Mean Absolute Error (MAE) for each coverage level, treating each level similar to a separate regression task. To maintain simplicity, we include the details about MAE in our supplementary materials.
@@
In the context of on-board data management proposed in Fig. 2, the false alarm ratios due to misclassifications indicate either the potential loss of data or the transmission of irrelevant data during downlink, as explained next. In cloud segmentation, the True Positive Ratio (TPR) involves correctly detecting cloudy pixels, while the True Negative Ratio (TNR) relates to recognizing non-cloudy data with potential value. The False Positive Ratio (FPR) is crucial because it indicates misclassified pixels, wrongly identified as clouds but that actually correspond to sea or land data. Following Fig. 2, these missclassifications can result in the discard or selective lossy compression with lower bitrate of valuable data. On the other hand, the False Negative Ratio (FNR) is less critical than FPR, as it does not involve data loss or data degradation, but only a less efficient usage of the downlink channel transmitting undetected pixels actually corresponding to clouds. In the context of sea segmentation, true positives signify correct identification of valuable sea data, while true negatives correctly recognize non-relevant data for potential deletion or compression. In sea-focused campaigns, FNR is critical, as it deals with misclassified sea-related pixels, potentially leading to loss of valuable data or degradation. In contrast, FPR is of less significance since its impact lies merely in downlink efficiency with the transmission of irrelevant pixels (clouds or land), rather than causing any actual data loss. The same analysis and conclusion apply when examining the context of land segmentation. In short, without taking into account the earlier suggested reasonable compromise between false positives and false negatives, for both sea and land categories, FNR has priority over FPR due to valuable data loss or degradation in the case of high FNR. However, for cloud data, FPR is critical as it leads to irreversible data loss or lossy compression, with FNR affecting merely downlink efficiency.
@@
(a) Segmentation by 1D-Justo-LiuNet.
(b) Segmentation by 2D-CUNet++ Reduced.
Figure 6: Segmentation results for a HYPSO-1 capture over the Namib desert in Namibia, Africa, close to the Gobabeb Namib Research Institute for dry land research. The approximate coordinates for this location are 23°19’06.0”S latitude and 14°44’21.8”E longitude. The grayscale image in (a) and (b) corresponds to the same L1b channel with a wavelength $\lambda$ of 699.61 nm, close to NIR spectrum. The term SAMPLES denotes the swath width captured across the direction of satellite travel, while LINES indicates the number of scan frames collected along the satellite’s path. As stated in [5, 29], HYPSO-1’s spatial resolution in the across- and along-track directions is approximately 100 m x 600 m.
@@ 3 Results
We conduct model training using 112 channels of both raw sensor data and L1b calibrated radiance, and we discuss the segmentation of raw and L1b radiance in Section 4. For L1b radiance, we present the outcomes of the segmentation models in Table II, where we utilize bold font to accentuate our models. The table displays the models’ performance for segmentation, with the primary metric being the average accuracy across the sea, land, and cloud classes. Additionally, to assess the sea-land-cloud segmentation performance for the context of the proposed on-board ranking system, we also include the average Spearman’s rank coefficient. In particular, the average Spearman’s coefficient across the respective rankings based on sea, land, and cloud coverage levels is presented in Table II. Moreover, we also include information about the network’s parameter count to offer insights of the potential memory demands for in-orbit deployment. In addition, we provide the average inference time per HS image measured in our testing setup on ground, on the NVIDIA RTX A4000 GPU (for DL), and on the 13th Gen Intel(R) Core (TM) i9-13900K CPU (for ML). From the numbers presented in Table II, we sort the models in Table III based on their respective metrics. In Table III, we first sort the models solely based on their average accuracy, from highest to lowest, and subsequently, we sort them in the following columns, first by the Spearman’s coefficient (highest to lowest), then by the number of parameters (lowest to highest), and finally by inference time (shorter to longer). Therefore, the first rows in the table represent the top models in the test set for each respective metric. However, while average accuracy offers a general overview on segmentation performance across the sea, land, and cloud classes, it may obscure poor model performance in specific classes. To address this, we offer additional insights into individual class performance by presenting their respective binary false detection ratios. Namely, in Figs. 3, 4, and 5, we provide a comparative visual representation of false alarm ratios for the binary segmentation of each distinct class. We note that the top-performing models get positioned in the lower-left corners, indicating the optimal trade-off balance between low false positives and low false negatives. Within the context of the example case presented in Fig. 2, the graphs illustrate the false alarm ratio on the Y-axis corresponding to the loss (or degradation) of valuable data, while the X-axis pertains to a false detection ratio that reflects the inefficient utilization of the downlink channel. In relation to this trade-off, we stress that while minimizing the loss of valuable data remains crucial, we also highlight the importance of maintaining an acceptable downlink efficiency.
Finally, our experiments conclude by training again the models using the 3-channel dataset instead. We employ the dataset to assess the impact of channel reduction on the segmentation models’ performance, parameter count, and inference times, as shown in Table IV. We note that the table omits the 1D models due to the 3-channel dataset’s restricted spectral information, as mentioned earlier in Section 2.3.4. Moreover, the empirical results are further confirmed by training once again the models with a set of three channels different from the ones in Section 2.2/A. The new results align with the previous findings and for simplicity, we include the new figures in the supplementary material.
Finally, as previously noted, the dataset predominantly contains instances of sea and land, with clouds representing the minority class. Consequently, the test set also exhibits a lower proportion of cloud data (14.13%), in contrast to sea (39.01%) and land (46.85%). We recognize this as a limitation that demands future refinement. Therefore, to mitigate the impact of the class imbalance during the testing phase, we conduct additional experiments. Namely, we perform inference on an additional on-ground deployment set, with 30 HS images, specifically designed to test the model’s generalization on e.g. cloudy and overexposed images to further support the findings of our work, and we provide the complete results in the supplementary material. As an example, Fig. 6 provides a comparison of the segmentation outcomes of the 1D-Justo-LiuNet and the 2D-CUNet++ Reduced for a sample image from the deployment set.
@@ 4 Discussion
4.1 Results Analysis for Models Trained on 112 Channels
We first analyze the results for the models trained on 112 channels (Tables II and LABEL:Table:models_sorted_according_to_performance). Our objective is to investigate how the utilization of the imaging spatial context by 2D models affects the sea-land-cloud segmentation across the test images. We assess the models’ strengths and limitations in terms of performance, spanning from the average accuracy and the binary false alarm ratios to the Spearman’s correlation coefficient for the ranking system. We finalise the analysis by discussing the model’s parameter count and inference time, which are relevant metrics for on-board deployment. Prior to the analysis of the results, we intuitively formulate a hypothesis next, independent of the empirical findings. The hypothesis asserts that it is reasonable to anticipate better performance for 2D models compared to 1D models. This expectation arises because 2D models, unlike 1D models relying solely on spectral context, analyse a richer context comprising both spectral and spatial information. We next proceed to elaborate whether the empirical results substantiate such hypothesis.
@@ 4.1.1 Segmentation Performance (I): Average Accuracy
When examining the segmentation performance solely based on the average accuracy across sea, land, and cloud classes, we observe that our proposed model, 1D-Justo-LiuNet, achieves the highest accuracy of 0.93 in Table II. The result of a 1D model outperforming 2D models is not an exception, as four out of five models ranking among the top performers in accuracy in Table LABEL:Table:models_sorted_according_to_performance are consistently 1D models. The only exception in this top five ranking is our 2D model, 2D-Justo-UNet-Simple, which achieves an accuracy of 0.92, closely approaching the performance of the 1D-Justo-LiuNet. Therefore, the findings lead to the conclusion that the initially expected hypothesis suggesting that 2D models would outperform 1D models does not hold valid in our experiments, as the 2D models exhibit worse performance. This phenomenon can be attributed to the fact that 1D models, unlike 2D using the spatial context, are not affected by the modest spatial resolution of HYPSO-1’s images while not suffering from the spatial variations that may be practically irrelevant and misleading for pixel-level segmentation. Instead, 1D models dedicate their efforts to analyzing exclusively the spectral context, which is, certainly, a powerful domain in HS imagery. Additionally, Table LABEL:Table:models_sorted_according_to_performance demonstrates that our proposed models (1D-Justo-LiuNet, 2D-Justo-UNet-Simple, 1D-Justo-HuNet, and 1D-Justo-LucasCNN) outperform, in average accuracy, the state-of-the-art models under analysis, when trained on the HYPSO-1 dataset for sea-land-cloud segmentation. Indeed, we surpass models such as the 2D-NU-Net-mod, 2D-NU-Net, 2D-UNet FAUBAI, and 2D-UNet FAUBAI Reduced previously employed in the field of cloud segmentation for satellite remote sensing imagery. Finally, we highlight that the average accuracy of 1D-Justo-LiuNet (0.93), 1D-Justo-HuNet (0.91), and 1D-Justo-LucasCNN (0.91) also surpass the literature’s average accuracy for soil texture [45]. As a matter of fact, the state of the art achieves significantly lower average accuracy of 0.59, 0.61, and 0.56 for Liuetal, Huetal, and LucasCNN, respectively.
@@ 4.1.2 Segmentation Performance (II): The Effect of L1b Radiance Calibration
The models in Table II are trained on L1b calibrated radiance data. L1b radiance is one of the standard data formats in HS imagery, which not only allows satellite comparisons but also removes sensor-specific characteristics and biases. Next, we empirically examine how L1b radiance calibration impacts segmentation performance. In the supplementary materials of this article, we provide the average accuracy, along with other performance metrics, for the models trained on raw sensor data. This supplementary information is provided to facilitate comparisons with the models trained on L1b radiance, as presented in Table II, in support of the upcoming analysis. To begin, for the highest-performing DL models in terms of average accuracy shown in Table III (1D-CNNs), the models exhibit slight superior performance with L1b radiance compared to training them with raw data. A similar trend occurs for the 2D models ranked just below the 1D models, which also demonstrate enhanced performance with radiance data versus raw data. Conversely, the models at the lower end of the accuracy ranking tend to perform better with raw data than radiance calibration. Concentrating only on the superior models in Table III, which hold more significance as they are the top performers, we conclude that L1b radiance calibration tends to positively impact and improve segmentation performance. We attribute the improvement to the correction of errors and biases during L1b radiance calibration, allowing top-performing models to focus on inherent data patterns rather than overfitting on sensor-induced characteristics, leading to improved generalization performance. Nonetheless, we strongly recommend conducting comparative testing on both calibrated and raw sensor data when training the models on datasets distinct from this paper to verify whether calibrated data, such as radiance or reflectance, continues to enhance model performance.
@@
4.1.3 Segmentation Performance (III): False Detection Ratios, Data Loss, and Downlink Inefficiency
@@
For a more in-depth assessment of the segmentation performance for each individual class, we scrutinize which models are positioned in the lower-left corner in the scatter plots in Figs. 3, 4, and 5. The lower-left region ensures the most optimal false detection ratios, and for the example ranking system in Fig. 2, the region achieves minimal loss (or degradation) of valuable data, while also minimizing the inefficient utilization of the satellite downlink. As previously explained, lower data loss (in the Y-axis) takes precedence, while the downlink inefficiency (in the X-axis) is of lower priority since it does not result in any loss of valuable data. However, as discussed in Section 2.4, we also aim for a reasonably balanced relationship between false positives and false negatives to prevent scenarios where a model achieves a convenient minimal data loss, at the expense of extremely deteriorating the downlink efficiency. Next, we comparatively analyze the top-performing models located in the lower-left corners of Figs. 3-5.
@@
First, we assess the models only in terms of data loss. In Fig. 3 for cloud segmentation, models such as the 2D-Justo-UNet-Simple excels in minimizing the loss of data, approaching nearly no data loss (ideal case). This signifies that the 2D-Justo-UNet-Simple outperforms by not misclassifying valuable non-cloud data (sea or land) as clouds, even if it means some cloudy data remains undetected, thus reducing the downlink efficiency. With respect to sea segmentation (Fig. 4), the 2D-UNet-FAUBAI-reduced stands out for its minimal loss of sea pixels with FNR=0.01, i.e., there is only a 1% proportion of sea pixels that remain undetected, resulting in their loss. This outperforms other models such as the 1D-Justo-LiuNet with 5% loss. Finally, in Fig. 5 for land segmentation, we observe a remarkable similarity in false alarm ratios in the lower-left optimal region, resulting in a slight overlap. Therefore, we next clarify that the optimal models, when sorted in ascending order of FNR (data loss), are as follows: 0.02 (2D-Justo-UNet-Simple), 0.03 (1D-Justo-LiuNet), 0.04 (1D-Justo-HuNet), and 0.06 (2D-CUNet++ Reduced). Consequently, the 2D-Justo-UNet-Simple is the most optimal at minimizing the loss of land data, achieving only a 2% ratio of undetected land pixels.
@@
Second, to jointly maintain a reasonable downlink efficiency while achieving low data loss, we study the models taking into account not only the data loss as in the previous analysis, but also the trade-off between both loss of data and downlink efficiency objectives. We could consider the use of Pareto fronts, which are typically applied in multi-objective optimization to address trade-offs between inversely proportional variables. Nevertheless, for simplicity, we analyze the trade-off using the Euclidean distance metric that gives equal weight to the two objectives. Specifically, we calculate the distance of each model from the origin point, representing the ideal scenario with zero false alarms. First, for cloud segmentation (Fig. 3), the models closest to the ideal performance have (FPR, FNR) values of (0.02, 0.15) for 1D-Justo-LiuNet and (0.00, 0.17) for 2D-Justo-UNet-Simple, resulting in approximate distances of 0.151 and 0.170, respectively. In brief, 1D-Justo-LiuNet exhibits the shortest distance to the ideal model. Second, the top models for sea segmentation (Fig. 4), with (FNR, FPR) of (0.08, 0.01) for 1D-Justo-LucasCNN, (0.05, 0.02) for 1D-Justo-LiuNet, and (0.04, 0.03) for 2D-UNet-FAUBAI, have approximate distances of 0.081, 0.054, and 0.050, respectively. This suggests a marginally better performance of 2D-UNet-FAUBAI, with a difference of 0.004 in distance compared to the 1D-Justo-LiuNet. However, Table II notes that the 2D-UNet-FAUBAI has over 26 million parameters, whereas our network, 1D-Justo-LiuNet, has only 4,563 parameters, which is more suitable for on-board deployment. Considering the nearly identical performance and the large parameter difference against FAUBAI’s network, 1D-Justo-LiuNet remains the top model also for sea segmentation. Finally, for land segmentation (Fig. 5), the optimal models are of (FNR, FPR) = (0.02, 0.06) for 2D-Justo-UNet-Simple, (0.03, 0.04) for 1D-Justo-LiuNet, (0.04, 0.04) for 1D-Justo-HuNet, and (0.06, 0.04) for 2D-CUNet++ Reduced, with approximate distances of 0.063, 0.050, 0.057, and 0.072, respectively. In summary, 1D-Justo-LiuNet outperforms the other models for land segmentation.
@@
Considering the initial analysis focused solely on data loss and the subsequent analysis regarding the trade-off between both data loss and downlink efficiency, we next provide our conclusion. On the one hand, the 2D-Justo-UNet-Simple, with approximately 7 thousand parameters, minimizes the loss of valuable data for most classes. The only exception is the sea class, where it incurs an acceptable 5% data loss in contrast to the 1% loss by 2D-UNet-FAUBAI-reduced which uses around 2 million parameters and hence the FAUBAI’s network is less suitable for on-board inference. On the other hand, when we equally consider both data loss and downlink usage efficiency, our model, 1D-Justo-LiuNet, excels in minimizing data loss, achieving low rates of 2%, 5%, and 3%, while simultaneously demonstrating relatively minimal inefficiency rates of 2%, 3%, and 15% for sea, land, and cloud classes, respectively. The findings suggest that when employing 1D-Justo-LiuNet, downlink efficiency slightly deteriorates due to an under-detection of clouds, leading to an overabundance of 15% of cloud data that would be downlinked from the satellite. We attribute the finding to the dataset’s class imbalance, where the cloud class constitutes the minority. Nevertheless, the over-detection of the majority classes sea (2%) and land (3%) pixels is smaller, thereby increasing the overall efficiency of the downlink channel.
4.1.4 Segmentation Performance (IV): Priority Rankings for Downlink